# Grassroots Platforms with Atomic Transactions: Social Networks, Cryptocurrencies, and Democratic Federations

### v2

@@ Abstract.
Grassroots platforms aim to offer an egalitarian alternative to global platforms — centralized/autocratic (Facebook etc.) and decentralized/plutocratic (Bitcoin etc.) alike. Whereas global platforms can have only a single instance–one Facebook, one Bitcoin—grassroots platforms can have multiple instances that emerge and operate independently of each other and of any global resource except the network, and can interoperate and coalesce into ever-larger instances once interconnected, potentially (but not necessarily) forming a single instance. Key grassroots platforms include grassroots social networks, grassroots cryptocurrencies, and grassroots democratic federations. Previously, grassroots platforms were defined formally and proven grassroots using unary distributed transition systems, in which each transition is carried out by a single agent. However, grassroots platforms cater for a more abstract specification using transactions carried out atomically by multiple agents, something that cannot be expressed by unary transition systems. As a result, their original specifications and proofs were unnecessarily cumbersome and opaque.
@@
Here, we aim to provide a more suitable formal foundation for grassroots platforms. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notion of grassroots platforms within this new foundation. We present crisp specifications of key grassroots platforms using atomic transactions: befriending and defriending for grassroots social networks, coin swaps for grassroots cryptocurrencies, and communities forming, joining, and leaving a federation for grassroots democratic federations. We prove a general theorem that a platform specified by atomic transactions that are so-called interactive is grassroots; show that the atomic transactions used to specify all three platforms are interactive; and conclude that the platforms thus specified are indeed grassroots. We thus provide a better mathematical foundation for grassroots platforms and a solid and clear starting point from which their implementation can commence.
@@
Grassroots Systems, Distributed Transition Systems, Atomic Transactions, Social Neworks, Cryptocurrencies, Democratic Online Communities
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Peer-to-peer architecturesccs: Networks Network protocol designccs: Networks Formal specificationsccs: Software and its engineering Distributed systems organizing principles
1. Introduction
@@
Background. The Internet today is dominated by centralised global platforms—social networks, Internet commerce, ‘sharing-economy’—with autocratic control (Zuboff, 2019, 2022). An alternative is emerging — blockchains and cryptocurrencies (Ethereum, ndao; Faqir-Rhazoui et al., 2021; Nakamoto and Bitcoin, 2008a; Wood et al., 2014; Wang et al., 2018, 2021)— that are also global platforms, but with decentralized, plutocratic control (Buterin, 2018).
@@
Grassroots platforms (Shapiro, 2023a, b, 2024; Halpern et al., 2024) aim offer an egalitarian alternative to global platforms, centralized and decentralized alike. Global platforms can only have a single instance–one Facebook, one Bitcoin—as two instances of the platform would clash over the same domain name/port number/boot nodes, which are hardwired into their code. It is possible to modify their code (‘hard-fork’) to create non-conflicting instances—Facebook, Bitcoin—which would ignore, rather than interoperate with, the primary instance. Grassroots platforms, in contrast, can have multiple instances that emerge and operate independently of each other and of any global resource except the network, and can interoperate and coalesce once interconnected, forming ever-larger platform instances, potentially (but not necessarily) coalescing into a single instance. Any platform that operates on a shared global resource or employs a single replicated (Blockchain (Nakamoto and Bitcoin, 2008b)), or distributed (IPFS (Benet, 2014), DHT (Rhea et al., 2005)) shared global data structure, and distributed pub/sub systems with a global directory (Chockler et al., 2007a, b; Buchegger et al., 2009)), are all not grassroots. An example of a platform that is grassroots in spirit (even if not formally proven as such) is Scuttlebutt (Tarr et al., 2019; Kermarrec et al., 2020). BitTorrent (BitTorrent, 2025) is similar in spirit but is peer-to-peer among servers not people, who may connect to a server from a personal device without a fixed IP address that does not participate in the file-dissemination protocol.
@@
Motivation. Key grassroots platforms include grassroots social networks (Shapiro, 2023b), grassroots cryptocurrencies (Shapiro, 2024; Lewis-Pye et al., 2023), and grassroots democratic federations (Halpern et al., 2024). Previously, grassroots platforms were defined formally using unary distributed transition systems, in which each transition is carried out by a single agent (Shapiro, 2023a). However, grassroots platforms cater for a more abstract specification using transactions carried out atomically by multiple agents, something that cannot be expressed by unary transition systems. As a result, their original specifications were more cumbersome and opaque than they should be. Moreover, in grassroots platforms transactions carried out by different sets of participants need not be synchronized with each other, which greatly simplifies their mathematical foundations.
@@
Here, we aim to provide a more suitable formal foundation for these key grassroots platforms and beyond. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notions of grassroots platforms and their implementation within this enhanced foundation.
@@
Furthermore, previous work (Shapiro, 2023a) provided sufficient conditions for when a platform (formally, a protocol defined via a family of distributed transition systems) is grassroots. While going through the route presented in previous work is possible, the new foundations offer another, more direct and mathematically preferable way to prove that a platform, specified via atomic transactions, is grassroots. Here, we follow this more direct path.
@@
Grassroots Social Networks. As a concrete example, consider grassroots social networks. Their goal is to provide people with the functionality of, say, Twitter/X-like feeds and WhatsApp-like groups, but without subjecting people, their social graph, or their personal information to the control of and exploitation by Musk, Zuckerberg, et al. A key component in a social network is the social graph, encoding who is a friend or a follower of whom. In centralized social networks (e.g. Twitter/Facebook/WhatsApp) the social graph is stored, controlled, and commercially-exploited (Zuboff, 2019, 2022) by the central authority. In proposed decentralized architectures for social networks (DSNP, porg), the social graph is stored on the globally-replicated blockchain, under the control of a third-party—the miners/validators that execute the blockchain consensus protocol, who are remunerated for their service.
@@
In a grassroots social network, the social graph is stored in a distributed way under the control of the people that participate in the network, with each person storing the local neighbourhood of the graph pertaining to them. Here is a fragment of the specification of the social-graph maintenance protocol of a grassroots social networks. It assumes that each agent maintains, as its local state, a set of friends; that two agents $p$ and $q$ can atomically become friends, if they were not friends beforehand; and that the two friends $p$ and $q$ can atomically cease to be friends. It employs two binary atomic transactions:
(1) Befriend: $c^{\prime}_{p}:=c_{p}\cup\{q\}$, $c^{\prime}_{q}:=c_{q}\cup\{p\}$, provided $q\notin c_{p}$ and $p\notin c_{q}$ (2) Defreind: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
Communication aspects of a social network can be added easily, under the restriction that communication occurs only among friends (Shapiro, 2023b). For example, feeds with followers, where friends follow share feeds with each other, and if two friends follow a third person, then the first to obtain an item on the third person’s feed disseminates it to the other. Formally, a liveness theorem can be proven for this design (Shapiro, 2023b), stating that if a person $p$ follows a person $q$ and is connected to $q$ via a chain of mutual friends, each of them correct and follows $q$, then $p$ will eventually receive every item on $q$’s feed. Practically, a true “network celebrity” employing this protocol can easily achieve efficient large-scale distribution of their feed via the friendship subgraph of their followers.
@@
Grassroots Cryptocurrencies. Grassroots cryptocurrencies (Shapiro, 2024; Lewis-Pye et al., 2023) can provide a foundation for an egalitarian digital economy, where people price their goods and services in terms of their own personally-minted grassroots coins, and liquidity is achieved through the creation of mutual credit lines among persons (natural and legal) via the swap of personal coins. Grassroots coins issued by different persons form an integrated digital economy via the sole rule in the grassroots cryptocurrencies protocol:
Coin Redemption: If $p$ holds a $q$-coin $c$ and $q$ holds any grassroots coin $c^{\prime}$, then $q$ must agree to swap $c$ for $c^{\prime}$, one-for-one, if so requested by $p$.
@@
(1) Mint: $c^{\prime}_{p}:=c_{p}\cup$¢${}_{p}^{k}$, $k>0$. (2) Swap: $c^{\prime}_{p}:=(c_{p}\cup y)\setminus x$, $c^{\prime}_{q}:=(c_{q}\cup x)\setminus y$, provided $x\subseteq c_{p}$ and $y\subseteq c_{q}$.
@@
  2. (2)
    Mutual Credit Lines: Establish mutual credit lines via the swap of personal (self-minted) coins, namely $x=$¢${}_{p}^{j}$ and $y=$¢${}_{q}^{k}$ for some $j,k>0$. Typically $j=k$, but mutual credit with a premium for one of the agents with $j\neq k$ is also possible. A digital economy based on grassroots cryptocurrencies achieves liquidity through persons (natural and legal) establishing mutual credit lines among them.
  3. (3)
@@
Grassroots Democratic Federation. Grassroots Democratic Federation aims to address the democratic governance of large-scale decentralized digital communities, e.g., of the size of the EU, the US, existing social networks, and even humanity at large. A grassroots democratic federation evolves via the grassroots formation and federation of digital communities, each governed by an assembly selected by sortition from its members. The approach assumes digital freedom of assembly—that people can freely form digital communities that can federate into ever-larger communities as well as spawn child communities, based on geography, relations, interests, or causes. Each community, no matter how large, is governed democratically by a small assembly elected by sortition among its members.
@@
The fundamental difference between grassroots social networks and grassroots federations is that the actors in the former are individuals, whereas the actors in the latter are communities of individuals, governed by assemblies. The standard centralized realization of this requirement would be to run a program on a server/cloud for each community, which realizes some democratic governance protocol by the assembly of that community. The standard decentralized realization of this requirement would be to similarly run a smart contract (De Filippi et al., 2021), realizing a Decentralized Autonomous Organization (DAO) (Ethereum, ndao) for each community on some decentralized global platform that providers a consensus protocol needed to execute smart contracts. Neither is grassroots. The grassroots solution would be to realize each community as a digital social contract (Cardelli et al., 2020), which also requires a consensus protocol among the participants, but one that is executed in a grassroots way by the participants themselves, not by third-parties (miners, validators) operating a global platform.
A specification of grassroots federation via unary distributed transition systems has not been attempted, but if it were, it would require specifying a consensus protocol by which members of two communities decide, atomically, to join or leave. Here, we define such atomic transactions abstractly: All members of a community carry its state locally; and all of them change its state atomically to realize joining or leaving another community. The details are shown below.
Paper outline. Section 2 introduces distributed transition systems and how they are induced by a set of atomic transactions. Section 3 provides transactions-based specifications of three key grassroots platforms: Grassroots social networks, grassroots cryptocurrencies, and grassroots democratic federations. Section 4 introduce protocols, grassroots protocols, protocols induced by atomic transactions, interactive atomic transactions, and proves the main theorem: That a protocol induced by an interactive set of transactions is grassroots. It then shows that the key platforms are interactive, and conclude that they are grassroots as specified. Section 5 discusses, informally, the implementation of grassroots platforms. Section 6 discussed related and future work.
2. Distributed Transition Systems and Grassroots Protocols
@@
2.1. Transition systems
The following is a simplified variation, sufficient for the purpose of this work, on the foundations introduced in (Shapiro, 2021). In the following, $a\neq b\in X$ is a shorthand for $a\neq b\wedge a\in X\wedge b\in X$.
@@ 2.1. Transition systems
We note that the definition of transition systems originally employed in the definition of grassroots systems (Shapiro, 2021, 2023a) included a liveness condition. As the rather-abstract atomic transactions employed here are volitional and have no associated liveness conditions, these, are omitted to simplify the exposition. Also, the original work considered faulty computations and fault-tolerant implementations, not considered here. The simplified-away components could be re-introduced in follow-on work that considers live and fault-tolerance implementations of the specification presented here.
2.2. Distributed Transition Systems with Atomic Transactions
@@
Unary distributed transition systems were introduced in (Shapiro, 2021) and were employed to define the notion of grassroots protocols (Shapiro, 2023a) and to provide unary specifications for various grassroots platforms (Shapiro, 2023b, 2024; Lewis-Pye et al., 2023). Here, we employ $k$-ary transition systems, for any $k\leq|P|$, in which several agents can change their state simultaneously. However, rather than specifying a transition system directly, we specify it via the atomic transactions (Lynch et al., 1988) the transition system intends to realize.
@@
  2. (2)
    Specifies explicitly its participants (both active and stationary), thus implicitly defining infinitely-many distributed transitions where non-participants remain in their arbitrary state.
In addition to being atomic, the description above implies that transactions are asynchronous (Shapiro, 2021), in that if a transaction can be carried out by its participants, with non-participants being in any arbitrary states, it can still be carried out no matter what the non-participants do. In particular, the active participants in a transaction need not synchronize its execution with any non-participant (but they must synchronize, of course, with stationary participants, as changing the local state of any participant would disable the transaction). Informally, a transaction is but a distributed transition over its participants, which defines a set of distributed transitions over any larger set of agents. Formally:
@@
Note that while the set of distributed transitions $R{\uparrow}P$ in general may be over a larger set of agents than $R$, $T$ and $R$ are of the same degree, by construction.
Note that while each of the distributed transitions of a distributed transition system over $P$ is also over $P$, in a set of transactions each member is typically over a different set of participants. Thus, a set of transactions $R$ over $S$, each with participants in some $Q\subseteq P$, defines a distributed transition system over $S$ and $P$ as follows:
@@
3. Grassroots Platforms with Atomic Transactions

### v3

@@ Abstract
Grassroots platforms aim to offer an egalitarian alternative to global platforms — centralized/autocratic (Facebook etc.) and decentralized/plutocratic (Bitcoin etc.) alike. Whereas global platforms can have only a single instance—one Facebook, one Bitcoin—grassroots platforms can have multiple instances that emerge and operate independently of each other and of any global resource except the network, and can interoperate and coalesce into ever-larger instances once interconnected, potentially (but not necessarily) forming a single instance. Key grassroots platforms include grassroots social networks, grassroots cryptocurrencies, and grassroots democratic federations. Previously, grassroots platforms were defined formally and proven grassroots using unary distributed transition systems, in which each transition is carried out by a single agent. However, grassroots platforms cater for a more abstract specification using transactions carried out atomically by multiple agents, something that cannot be expressed by unary transition systems. As a result, their original specifications and proofs were unnecessarily cumbersome and opaque.
@@
Here, we aim to provide a more suitable formal foundation for grassroots platforms. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notion of grassroots platforms within this new foundation. We present crisp specifications of key grassroots platforms using atomic transactions: befriending and defriending for grassroots social networks, coin swaps for grassroots cryptocurrencies, and communities forming, joining, and leaving a federation for grassroots democratic federations. We prove a general theorem that a platform specified by atomic transactions that are so-called interactive is grassroots; show that the atomic transactions used to specify all three platforms are interactive; and conclude that the platforms thus specified are indeed grassroots. We thus provide a crisp mathematical foundation for grassroots platforms and a solid and clear starting point from which their implementation can commence.
keywords:
@@
Grassroots Systems, Distributed Transition Systems, Atomic Transactions, Social Networks, Cryptocurrencies, Democratic Online Communities
1 Introduction
@@
Background. The Internet today is dominated by centralised global platforms—social networks, Internet commerce, ‘sharing-economy’—with autocratic control [52, 53]. An alternative is emerging—blockchains and cryptocurrencies [23, 24, 35, 51, 48, 49]—that are also global platforms, but with decentralized, plutocratic control [12].
@@
Grassroots platforms [43, 44, 45, 25] aim offer an egalitarian alternative to global platforms, centralized and decentralized alike. Global platforms can only have a single instance–one Facebook, one Bitcoin—as two instances of the platform would clash over the global resources they utilize—domain name, port number, or boot nodes, which are hardwired into their code. Whether it is possible to modify their code (‘hard-fork’) to create non-conflicting instances—Facebook, Bitcoin—such forked instances would ignore, rather than interoperate with, the primary instance. Grassroots platforms, in contrast, can have multiple instances that emerge and operate independently of each other and of any global resource except the network, yet can interoperate and coalesce once interconnected, forming ever-larger platform instances, potentially (but not necessarily) coalescing into a single instance.
Any platform that operates on a shared global resource or employs a single replicated (Blockchain [36]), or distributed (IPFS [6], DHT [39]) shared global data structure, and distributed pub/sub systems with a global directory [15, 16, 10], are all not grassroots. BitTorrent [7] and Mastodon [38] are similar in spirit but are peer-to-peer among servers not people; in particular, in such systems a group of people with smartphones cannot do more if more people with smartphones joint the group, unless they are also joined by a server, thus these systems do not fall under the definition of grassroots. An example of a platform that is grassroots in spirit (even if not formally proven as such) is Scuttlebutt [47, 26].
@@
Motivation. Key grassroots platforms include grassroots social networks [44], grassroots cryptocurrencies [45, 29], and grassroots democratic federations [25]. Previously, grassroots platforms were defined formally using unary distributed transition systems, in which each transition is carried out by a single agent [43]. However, grassroots platforms cater for a more abstract specification using transactions carried out atomically by multiple agents, something that cannot be expressed by unary transition systems. As a result, their original specifications were more cumbersome and opaque than they should be. Moreover, in grassroots platforms transactions carried out by different sets of participants need not be synchronized with each other, which greatly simplifies their mathematical foundations.
@@
Here, we aim to provide a crisp formal foundation for these key grassroots platforms and beyond. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notions of grassroots platforms and their implementation within this enhanced foundation.
@@
Furthermore, previous work [43] provided sufficient conditions for when a platform (formally, a protocol defined via a family of distributed transition systems) is grassroots. While going through the route presented in previous work is possible, the new foundations offer a simpler, more direct and mathematically preferable way to prove that a platform, specified via atomic transactions, is grassroots. Here, we follow this simpler path.
@@
Grassroots Social Networks. As a concrete example, consider grassroots social networks. Their goal is to provide people with the functionality of, say, Twitter/X-like feeds and WhatsApp-like groups, but without subjecting people, their social graph, or their personal information to the control of and exploitation by Musk, Zuckerberg, et al. A key component in a social network is the social graph, encoding who is a friend or a follower of whom. In centralized social networks (e.g. Twitter/Facebook/WhatsApp) the social graph is stored, controlled, and commercially-exploited [52, 53] by the central authority. In proposed decentralized architectures for social networks [21], the social graph is stored on the globally-replicated blockchain, under the control of a third-party—the miners/validators that execute the blockchain consensus protocol, who are remunerated for their service.
@@
In a grassroots social network, the social graph is stored in a distributed way under the control of the people that participate in the network, with each person storing the local neighbourhood of the graph pertaining to them, and no third-party having access to any fragment of the social graph unless explicitly granted. Here is a fragment of the specification of the social-graph maintenance protocol of a grassroots social networks. It assumes that each agent maintains, as its local state, a set of friends; that two agents $p$ and $q$ can atomically become friends, if they were not friends beforehand; and that the two friends $p$ and $q$ can atomically cease to be friends. It employs two binary atomic transactions:
1. Befriend: $c^{\prime}_{p}:=c_{p}\cup\{q\}$, $c^{\prime}_{q}:=c_{q}\cup\{p\}$, provided $q\notin c_{p}$ and $p\notin c_{q}$ 2. Defreind: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
Communication functions of a social network can be added, under the restriction that communication occurs only among friends [44]. For example, feeds with followers, where friends follow each other, and if two friends follow a third person, then the first to obtain an item on the third person’s feed disseminates it to the other. Formally, a liveness theorem can be proven for this design [44], stating that if a person $p$ that follows a person $q$ is connected to $q$ via a chain of mutual friends, each of them correct and follows $q$, then $p$ will eventually receive every item on $q$’s feed. Practically, a true “network celebrity” employing this protocol may achieve efficient large-scale distribution of their feed via the friendship subgraph of their followers.
@@
Grassroots Cryptocurrencies. Grassroots cryptocurrencies [45, 29] can provide a foundation for an egalitarian digital economy, where people price their goods and services in terms of their own personally-minted grassroots coins, and liquidity is achieved through the creation of mutual credit lines among persons (natural and legal) via the swap of personal coins. Grassroots coins issued by different persons form an integrated digital economy via the sole rule in the grassroots cryptocurrencies protocol:
Coin Redemption: $p$ can redeem any $q$-coin it holds against any coin held by $q$, one-for-one.
@@
1. Mint: $c^{\prime}_{p}:=c_{p}\cup$¢${}_{p}^{k}$, $k>0$. 2. Swap: $c^{\prime}_{p}:=(c_{p}\cup y)\setminus x$, $c^{\prime}_{q}:=(c_{q}\cup x)\setminus y$, provided $x\subseteq c_{p}$ and $y\subseteq c_{q}$.
@@
  2. 2.
    Mutual Credit Lines: Establish mutual credit lines via the swap of personal (self-minted) coins, namely $x=$¢${}_{p}^{j}$ and $y=$¢${}_{q}^{k}$ for some $j,k>0$. Typically $j=k$, but mutual credit with a premium for one of the agents, namely with $j\neq k$, is also possible. A digital economy based on grassroots cryptocurrencies achieves liquidity through persons (natural and legal) establishing mutual credit lines among them.
  3. 3.
@@
Grassroots Democratic Federation. Grassroots Democratic Federation aims to address the democratic governance of large-scale decentralized digital communities, e.g., of the size of the EU, the US, existing social networks, and even humanity at large. A grassroots democratic federation evolves via the grassroots formation and federation of digital communities, each governed by an assembly selected by sortition from its members. The approach assumes digital freedom of assembly—that people can freely form digital communities that can federate into ever-larger communities as well as spawn child communities, based on geography, relations, interests, or causes. Small communities (say $<100$) are governed by their members. Larger communities—no matter how larger—are each governed democratically by a small (say $\approx 100$) assembly elected by sortition among its members.
@@
The fundamental difference between grassroots social networks and grassroots federations is that the actors in the former are individuals, whereas the actors in the latter are communities of individuals, governed by assemblies. The standard centralized realization of this requirement would be to run a program on a server/cloud for each community, which realizes some democratic governance protocol by the assembly of that community. The standard decentralized realization of this requirement would be to similarly run a smart contract [17], realizing a Decentralized Autonomous Organization (DAO) [23] for each community on some decentralized global platform that providers a consensus protocol needed to execute smart contracts. Neither is grassroots. The grassroots solution would be to realize each community as a digital social contract [13], which also requires a grassroots consensus protocol among the participants, namely one that is executed by the participants themselves, not by third-parties (miners, validators) operating a global platform.
A specification of grassroots federation via unary distributed transition systems has yet to be attempted. It would require a grassroots consensus protocol by which members of two communities decide, atomically, to join or leave. Here, we define such atomic transactions abstractly: All members of a community carry its state locally; and all of them change its state atomically to realize joining or leaving another community. The details are shown below.
Paper outline. Section 2 introduces distributed transition systems and how they are induced by a set of atomic transactions. Section 3 provides transactions-based specifications of three key grassroots platforms: Grassroots social networks, grassroots cryptocurrencies, and grassroots democratic federations. Section 4 introduce protocols, grassroots protocols, protocols induced by atomic transactions, interactive atomic transactions, and proves the main theorem: That a protocol induced by an interactive set of transactions is grassroots. It then shows that the three key platforms introduced here are interactive, and concludes that they are grassroots as specified. Section 5 discusses, informally, the implementation of grassroots platforms. Section 6 discussed related and future work.
2 Distributed Transition Systems and Grassroots Protocols
@@
2.1 Transition systems
The following is a simplified variation, sufficient for the purpose of this work, on the foundations introduced in [42]. In the following, $a\neq b\in X$ is a shorthand for $a\neq b\wedge a\in X\wedge b\in X$.
@@ 2.1 Transition systems
We note that the definition of transition systems originally employed in the definition of grassroots systems [42, 43] included a liveness condition. As the rather-abstract atomic transactions employed here are volitional and have no associated liveness conditions, these, are omitted to simplify the exposition. Also, the original work considered faulty computations and fault-tolerant implementations, not considered here. The simplified-away components could be re-introduced in follow-on work that considers live and fault-tolerance implementations of the specification presented here.
2.2 Distributed Transition Systems with Atomic Transactions
@@
Unary distributed transition systems were introduced in [42] and were employed to define the notion of grassroots protocols [43] and to provide unary specifications for various grassroots platforms [44, 45, 29]. Here, we employ $k$-ary transition systems, for any $k\leq|P|$, in which several agents can change their state simultaneously. However, rather than specifying a transition system directly, we specify it via the atomic transactions [30] the transition system intends to realize.
@@
  2. 2.
    Specifies explicitly its participants (both active and stationary), thus implicitly defining infinitely-many distributed transitions where all non-participants remain in their arbitrary state.
In addition to being atomic, the description above implies that transactions are asynchronous [42], in the sense that if a transaction can be carried out by its participants, with non-participants being in any arbitrary states, it can still be carried out no matter what the non-participants do. In particular, the active participants in a transaction need not synchronize its execution with any non-participant (but they must synchronize, of course, with stationary participants, as changing the local state of any participant would disable the transaction). Informally, a transaction is but a distributed transition over its participants, which defines a set of distributed transitions over any larger set of agents. Formally:
@@
Note that while the set of distributed transitions $R{\uparrow}P$ in general may be over a larger set of agents than $R$, $T$ and $R$ are of the same degree, by construction. Also, note that while each of the distributed transitions of a distributed transition system over $P$ is also over $P$, in a set of transactions each member is typically over a different set of participants. Thus, a set of transactions $R$ over $S$, each with participants in some $Q\subseteq P$, defines a distributed transition system over $S$ and $P$ as follows:
@@
3 Grassroots Platforms with Atomic Transactions